OpenStudy (anonymous):

Find the exact value. Sec^2(11pi/6)

OpenStudy (jdoe0001):

hmmm kinda missing the \(\pi\) there.. shoot

OpenStudy (jdoe0001):

\(\bf sec(\theta) =\cfrac{1}{cos(\theta)} \\ \quad \\ sec^2\left( \frac{11\pi }{6}\right)\implies \left[ sec\left( \frac{11\pi }{6}\right) \right]^2\implies \left[ \cfrac{1}{cos\left( \frac{11\pi }{6}\right)} \right]^2 \\ \quad \\ \cfrac{1^2}{cos^2\left( \frac{11\pi }{6}\right)}\implies \cfrac{1}{cos^2\left( \frac{11\pi }{6}\right)}\implies \cfrac{1}{\left[cos\left( \frac{11\pi }{6}\right)\right]^2}\)

OpenStudy (jdoe0001):

so just get \(\bf cos\left( \frac{11\pi }{6}\right)\) from the Unit Circle, and square it

OpenStudy (anonymous):

Thank you so much!

OpenStudy (jdoe0001):

yw